# Empirically Validating Assumptions for Reliable A/B-Testing

### v1

@@ 1. Introduction & Motivation
In this work, we describe a practical approach to empirically validate whether we can trust the CIs and resulting $p$-values. When the null hypothesis holds true by design, we would expect a CI with confidence level $100\cdot(1-\alpha)\%$ to exclude zero in $\alpha\%$ of cases. Furthermore, the distribution of $p$-values should be expected to be uniform. We can emulate such situations by repeatedly resampling synthetic A/A-tests from the available data, leveraging the Kolmogorov-Smirnov test to ascertain whether the resulting $p$-value distribution is likely to be uniform. This provides an efficient and effective approach to empirically validate the assumptions that permeate common practice, allowing us to pinpoint problematic experimental designs where the false positive rate deviates from $\alpha$ (Kohavi and Chen, 2024). This, in turn, enforces statistical rigour and increases the trust we can place in A/B-testing estimands that do pass the test.
@@ 3. Experimental Results & Discussion
Aside from event frequency (i.e. $\mathbb{E}[Y]$), we might be interested in other moments of the outcome distribution $\mathsf{P}(Y)$. We consider the sample skewness of $Y$, and report it alongside histograms for four events on Figure 3. We visualise, in order of the legend: (i) the most frequent event, (ii) the rarest event for which normality cannot be rejected, (iii) an event with similar frequency to (iv) but low $D$-statistic, and (iv) the most frequent event with $p$-value ¡ $41e-$ 4. Sample skewness estimates are shown in the legend, suggesting that a higher skewness implies slower CLT convergence. Spearman’s $\rho\approx 0.43$ suggests a significant rank-correlation with the $D$-statistic, whilst confirming the independent informational value of the latter.

### v2

@@ 1. Introduction & Motivation
In this work, we describe a practical approach to empirically validate whether we can trust the CIs and resulting $p$-values. When the null hypothesis holds true by design, we expect a CI with confidence level $100\cdot(1-\alpha)\%$ to exclude zero in $\alpha\%$ of cases. Furthermore, the distribution of $p$-values should be expected to be uniform. We can emulate such situations by repeatedly resampling synthetic A/A-tests from the available data, leveraging the Kolmogorov-Smirnov test to ascertain whether the resulting $p$-value distribution is likely to be uniform (Kohavi et al., 2020, p. 188). This provides an efficient and effective approach to empirically validate the assumptions that permeate common practice, allowing us to pinpoint problematic experimental designs where the false positive rate deviates from $\alpha$ (Kohavi and Chen, 2024). This, in turn, enforces statistical rigour and increases the trust we can place in A/B-testing estimands that do pass the test.
@@ 3. Experimental Results & Discussion
Aside from event frequency (i.e. $\mathbb{E}[Y]$), we might be interested in other moments of the outcome distribution $\mathsf{P}(Y)$. Kohavi et al. (2014) and (Kohavi et al., 2022) discuss the skewness as an important diagnostic for CLT appropriateness. We consider the sample skewness of $Y$, and report it alongside histograms for four events in Figure 3. We visualise, in order of the legend: (i) the most frequent event, (ii) the rarest event for which normality cannot be rejected, (iii) an event with similar frequency to (iv) but low $D$-statistic, and (iv) the most frequent event with $p$-value ¡ $41e-$ 4. Sample skewness estimates are shown in the legend, suggesting that a higher skewness implies slower CLT convergence. Nevertheless, there is no monotonic relationship: Spearman’s $\rho\approx 0.43$ suggests a significant rank-correlation with the $D$-statistic, but confirms the independent informational value of the Kolmogorov-Smirnov test as a diagnostic tool.